# Accepted to 2024 European Conference on Computer Vision

### v2

@@ I Introduction
Computer vision has experienced significant success over the last years due to the advancement of Artificial Neural Networks such as Convolutions Neural Networks (CNNs) and Vision Transformers. Recently, the attraction of Spiking Neural Networks (SNNs) has been growing, due to their potential in real-time sensory processing, closer imitation of biological brain functionalities, and low power consumption. The advent of event-based cameras further accelerates this evolution, offering a promising path to harness the strengths of SNNs and address the challenges faced by traditional vision systems. This combination paves the way for more energy-efficient and adaptive solutions in computer vision [1].
@@
SNNs’ performance benefits from homeostasis, i.e. maintaining a stable neuron activity level, neither spiking too often nor staying too quiet [2, 3]. To achieve this, standard methods adjust parameters like fixed firing thresholds or synaptic weights. These strategies require fine-tuning, and may not be versatile across network architectures. The fixed thresholds in spiking neurons, as shown in Fig. 1, can be sensitive to initial conditions, leading networks to be overly active or dormant. The inflexibility in adapting to varying inputs makes neurons either too sensitive or unresponsive, affecting negatively the performance. It also compromises energy efficiency, with either high power usage from excessive firing or missed data from sparse firing.
To address the above challenges, we introduce the Asynchronous Bioplausible Neuron (ABN) function, which can be incorporated into the two most prominent bioplausible neuron models: the Spike Response Model (SRM) [4] and the Leaky Integrate-and-Fire (LIF) model [5]. ABN is comprised of three novel elements MG, TRG, and SE. In the MG component, we are the first to propose calculating gradients to quantify the movement of membrane potential, akin to the ‘rate of membrane potential change’ observed in biological neurons, which reflects how neurons dynamically adjust their potential in response to inputs [6]. This provides the direction of the movement necessary to maintain the neuron’s output firing. In TRG, we are the first to implement inertia in threshold movement using a sliding window to control bursts of spikes, offering network-independent threshold control. This mirrors the ‘adaptive threshold’ or ‘activity-dependent threshold modulation’ observed in biological neurons, where firing thresholds adapt based on historical activity to maintain stability [7]. Although the principle of the SE element is adopted from existing literature, we are the first to implement it using a sliding window mechanism and asynchronous event data, relating it to ‘neuronal efficiency’ or ‘firing efficiency,’ which measures how effectively a neuron converts electrical inputs into action potentials, a crucial factor in preventing response saturation [8].
We validate the effectiveness of our approach using two SNNs for object segmentation and image classification tasks under both normal and various degraded conditions, such as low light, high occlusion, etc. Further, our method is tested on six different datasets: N-MNIST [9], ESD-1, ESD-2 [10], DVS128-Gesture [11], N-ImageNet [12] and CIFAR-10 DVS [13] to demonstrate its general applicability. Additionally, we conducted an extensive study of the homeostasis (i.e. stability test of neuron firing), an ablation study to understand the impact of elements of the proposed function, and compared the power consumption with state-of-the-art methods.
@@ II-A Computer vision applications of SNNs
The exploration of SNNs for object segmentation and image classification has ignited widespread interest in the field, as the SNNs replicate the biological neuronal system to provide efficient and effective solutions for this task. Initial works in object segmentation leaned on feed-forward SNN architectures, utilizing the spike-timing-dependent plasticity (STDP) learning rule for unsupervised learning of features [14]. The Backpropagation Through Time (BPTT) method introduced in [15] for training SNNs, offers a more structured and supervised learning approach that significantly enhances feature extraction and classification accuracy, yet it is limited by its high computational complexity, substantial memory requirements, susceptibility to vanishing and exploding gradients, challenges in capturing long-term dependencies, and inefficiency in real-time learning applications. The introduction of Convolutional Spiking Neural Networks (CSNNs) marked a significant development in image classification. By incorporating convolutional layers into the network, these models could better handle the spatial complexities inherent in images, significantly enhancing segmentation performance. Concurrently, they maintained the temporal dynamics native to SNNs, enabling efficient processing of streaming data [16]. Recently, there has been a surge in the application of recurrent spiking neural networks (RSNNs) to image classification. These networks leverage feedback connections, enhancing the network’s ability to capture temporal dependencies in the data. Furthermore, their capacity to retain information over extended periods makes them particularly suited to segmentation tasks involving sequential or video data [17].
The memory-based methods such as those utilizing LSTM and RNN principles, highlighted by [18] and [19], offer advancements yet face limitations due to their application across entire networks, constraining individual neuron-based threshold modeling. Similarly, Spiking Neural Networks (SNNs) have been applied to tasks like segmentation and object detection with promising results as demonstrated by [20] and [21]; however, these models often draw from Artificial Neural Network (ANN) principles, potentially veering away from the bio-plausibility sought to closely mimic biological brain functions. These developments underscore the challenges in achieving nuanced modeling and bio-inspired design within neural network research, highlighting the need for approaches that balance technical efficacy with the intricacies of biological mechanisms. Generative networks [22] in event-based vision struggle due to the asynchronous and unpredictable nature of the data, leading to performance issues in scenarios with sporadic events. This challenge highlights the complexity of creating bio-plausible models that maintain stability and accuracy in the face of such data variability.
@@ II-B Dynamic Thresholding in SNNs
Research on SNNs has delved deeply into dynamic thresholding methods to bolster adaptability. For instance, while predictive coding emphasizes anticipating input patterns, it overlooks the spike timing significance, an aspect STDP-based adjustments address [23], [24]. The holistic approach of network activity regulation contrasts with the inherent refractory period modulation, focusing on preventing rapid neuron firing [25]. Furthermore, spiking models like SRM and LIF have advanced neuron dynamics representations [4, 26]. However, gaps remain, especially in dynamic threshold modulation. [27] introduced a method utilizing a scaling factor to adjust the threshold in response to increased input. Yet, since it was tailored specifically for incremental spike membrane potential, it faltered when faced with variations in membrane decays. On the other hand, [28] adopted exponential functions to dictate decay. This approach, however, resulted in fluctuating thresholding levels, failing to ensure a smooth threshold curve. Meanwhile, [29] ventured into a firing counter-based threshold adjustment. While innovative, it necessitated meticulous fine-tuning and exhaustive experimentation, leaving the determination of an optimal count mired in ambiguity.
Previous research may not addressed all the dynamics in the input signal, especially in highly complex event-based vision signals. In the event-based vision, SNNs face unique challenges: Neurons within the same layer exhibit varying event counts; some neurons experience more events than others. Consequently, applying uniformly the BDETT method [30], which utilizes average membrane potential of the neurons in a layer to predict the threshold of all the neurons in that layer, fails to reflect the individual dynamics of each neuron. This paper addresses notable research gaps. No prior studies have demonstrated that a bioinspired dynamic threshold can stabilize spike firing in semantic segmentation tasks. Additionally, these studies typically employ simplistic datasets such as NMNIST [9] and rarely evaluate system robustness under challenging conditions like motion blur or low light, vital for assessing the system’s homeostasis [31].
@@ III-A1 Event-based cameras
Event-based vision cameras are designed to register variations in logarithmic light intensities by recording individual pixel-level changes, referred to as events. These events formulate a continuous asynchronous event stream, which can be mathematically articulated as a series of ordered tuples. A tuple corresponds to an event $i$ and includes its spatial coordinates $(x_{i},y_{i})$ in the image frame, timestamps $t_{i}$, and polarity value $z_{i}$ [32]:
@@ III-A2 Leaky integrate-and-fire
The LIF model, as shown in Fig. 1, is characterized by its linear integration of inputs until a threshold $\Theta(t)$ is reached, causing its firing [5]. Mathematically, the sub-threshold dynamics of the membrane potential $V(t)$ of an LIF neuron can be described by the following differential equation:
@@
where $\tau_{m}$ is the membrane time constant that dictates the speed at which the neuron responds to inputs, $E_{L}$ is the leak potential or resting potential representing the equilibrium potential of the neuron in the absence of any input, $R_{m}$ is the membrane resistance, and $I(t)$ represents the total synaptic input to the neuron that might come from other neurons or external sources. When the membrane potential $V(t)$ reaches the threshold $\Theta(t)$, the neuron emits a spike and then resets its potential to a reset value $V_{\text{reset}}$, typically below $\Theta(t)$. Additionally, after spiking, the neuron enters a refractory period for a duration $T_{\text{ref}}$ during which it cannot fire again. The "leaky" aspect of the LIF model comes from the term $(V(t)-E_{L})$ in the differential equation. This term ensures that in the absence of any input, the neuron’s potential will gradually revert to the leak potential $E_{L}$.
@@ III-A3 Spike response model
In the SRM [4], incoming spike trains $s_{i}(t)$ are initially transformed into a spike response signal by convolving them with a spike response kernel $\varepsilon(\cdot)$, denoted as $(\varepsilon\ast s_{i})(t)$. Subsequently, this spike response signal is scaled by a synaptic weight $w_{i}$. The depolarization process involves summing up these scaled spike response signals across all incoming spikes, mathematically represented as $\Sigma_{i}w_{i}(\varepsilon\ast s_{i})(t)$. The SRM also incorporates a mechanism for hyperpolarization when a spike $s(t)$ is triggered, characterized by a refractory potential $(\zeta\ast s)(t)$, where $\zeta(\cdot)$ serves as the refractory kernel. This dual-component approach enables SRM to capture both excitatory and inhibitory dynamics in neuronal spike activity.
@@ III-A4 Asynchronous events to spike conversion
Traditionally, rate coding and temporal coding [33], [34] are the popular methods to convert frames, from a traditional CMOS, into spiking trains. In this study, similar to [35], each event is treated as a spike to avoid the conversation of the input signal into a spike train. This approach essentially uses a form of temporal coding, because the information is conveyed through the timing of the spikes. We have chosen to drop the polarity aspect of the events due to its high sensitivity, which can introduce noise and complicate the signal processing. Mathematically, E(t) is the event stream, $e=(x,y,t)$ is an individual event with coordinates $x$, y and timestamp t, and $S_{i}(t)$ is the spike train for the i-th neuron:
@@
Here, $\delta(t-e_{t})$ is the Dirac delta function, which represents a "spike" at time $t$.
@@ III-B Asynchronous Bioplausible Neuron
The equation models the dynamic adjustment of the neuron’s firing threshold at the next time step $t+1$ based on various factors. In this equation, $\Theta(t)$ is the existing threshold at time $t$, and $k_{1},k_{2},k_{3}$ are weights that determine the relative influence of each term. The term $k_{1}\cdot MG_{i}$ increases the threshold when the membrane potential is rapidly changing, preventing the neuron from firing too easily. Conversely, the term $-k_{2}\cdot TRG_{i}$ decreases the threshold if there has been a historical trend of rapid threshold increases. Lastly, the term $k_{3}\cdot SE_{i}$ takes into account the neuron’s recent spiking activity. Specifically, if the neuron is overly active, this term will make it more difficult for the neuron to fire in the future, thereby ensuring efficiency. Overall, this equation encapsulates a multi-faceted approach to dynamically adapting the neuron’s firing threshold.
@@ III-B1 Membrane Gradient (MG)
The rate of change of the membrane potential is a critical parameter for predicting a dynamic threshold in SNNs. It serves as an indicator of the neuron’s responsiveness to incoming spikes and plays an essential role in governing the neuron’s firing behavior. Positive correlations between dynamic thresholds and membrane potentials have been observed in several areas of diverse biological nervous systems. In [36], the average membrane potential is correlated with an average threshold to predict the dynamic threshold. Instead, in this work we correlate the rate of change of membrane potential with a threshold. A high rate of change often signifies that the neuron is receiving a burst of spikes, necessitating a higher threshold to maintain neuronal homeostasis. Conversely, a lower rate of change could indicate that the neuron is less active, allowing for a lower dynamic threshold to be sensitive to sparse inputs. Incorporating the rate of change of the membrane potential into the calculation of a dynamic threshold ensures that the SNNs can adapt to varying synaptic activities, thereby emulating the adaptability and energy efficiency observed in biological neurons.
The BDETT method presented in [30] calculates the dynamic threshold using the average of membrane potentials and thresholds. In a layer with varied neuronal activity, not all neurons are always active. Including inactive neurons in this calculation can pull the average value lower. Applying this reduced average threshold to highly active neurons may result in excessive spiking, leading to overly aggressive neuronal activity. In contrast, in this work, we adopt the neuron-specific derivative term, denoted as $\eta\cdot\frac{dv_{i}(t)}{dt}$, which captures the rate of change of membrane potential of neuron $i$ at time $t$, denoted by $v_{i}(t)$. To numerically compute the derivative, a finite difference approximation is employed. This involves the subtraction of the membrane potential at the previous timestep from the current potential, followed by division by the timestep size $\Delta t$. The scaling factor $\eta$, which is set to a value less than 1, reduces the rate of change to ensure that the rate of change of the threshold remains less than the rate of change of the potential. The explicit equation for the derivative term is given as:

### v3

@@ I Introduction
Computer vision has experienced significant success over the last years due to the advancement of Artificial Neural Networks such as Convolutions Neural Networks (CNNs [Kachole2020AHoneybees, sanket20163Tracker] and Vision Transformers [kachole2024bimodal, kachole2023bimodal, sanket20163Tracker]. Recently, the attraction of Spiking Neural Networks (SNNs) has been growing, due to their potential in real-time sensory processing, closer imitation of biological brain functionalities, and low power consumption. The advent of event-based cameras further accelerates this evolution, offering a promising path to harness the strengths of SNNs and address the challenges faced by traditional vision systems [kachole2024asynchronous, naeini2022event]. This combination paves the way for more energy-efficient and adaptive solutions in computer vision [Pozo2010UnravelingPlasticity].
@@
SNNs’ performance benefits from homeostasis, i.e. maintaining a stable neuron activity level, neither spiking too often nor staying too quiet [Cazalets2023AnCombination, Yeung2004SynapticModel]. To achieve this, standard methods adjust parameters like fixed firing thresholds or synaptic weights. These strategies require fine-tuning, and may not be versatile across network architectures. The fixed thresholds in spiking neurons, as shown in Fig. 1, can be sensitive to initial conditions, leading networks to be overly active or dormant. The inflexibility in adapting to varying inputs makes neurons either too sensitive or unresponsive, affecting negatively the performance. It also compromises energy efficiency, with either high power usage from excessive firing or missed data from sparse firing.
To address the above challenges, we introduce the Asynchronous Bioplausible Neuron (ABN) function, which can be incorporated into the two most prominent bioplausible neuron models: the Spike Response Model (SRM) [Gerstner1995TimeModels.] and the Leaky Integrate-and-Fire (LIF) model [Gerstner2002SpikingModels]. ABN is comprised of three novel elements MG, TRG, and SE. In the MG component, we are the first to propose calculating gradients to quantify the movement of membrane potential, akin to the ‘rate of membrane potential change’ observed in biological neurons, which reflects how neurons dynamically adjust their potential in response to inputs [sanes1993activity]. This provides the direction of the movement necessary to maintain the neuron’s output firing. In TRG, we are the first to implement inertia in threshold movement using a sliding window to control bursts of spikes, offering network-independent threshold control. This mirrors the ‘adaptive threshold’ or ‘activity-dependent threshold modulation’ observed in biological neurons, where firing thresholds adapt based on historical activity to maintain stability [lu2003bdnf]. Although the principle of the SE element is adopted from existing literature, we are the first to implement it using a sliding window mechanism and asynchronous event data, relating it to ‘neuronal efficiency’ or ‘firing efficiency,’ which measures how effectively a neuron converts electrical inputs into action potentials, a crucial factor in preventing response saturation [yu2017energy].
We validate the effectiveness of our approach using two SNNs for object segmentation and image classification tasks under both normal and various degraded conditions, such as low light, high occlusion, etc. Further, our method is tested on six different datasets: N-MNIST [Orchard2015ConvertingSaccades], ESD-1, ESD-2 [Huang2023AEnvironment, huang2024neuromorphic], DVS128-Gesture [Amir2017ASystem], N-ImageNet [kim2021n] and CIFAR-10 DVS [Li2017CIFAR10-DVS:Classification] to demonstrate its general applicability. Additionally, we conducted an extensive study of the homeostasis (i.e. stability test of neuron firing), an ablation study to understand the impact of elements of the proposed function, and compared the power consumption with state-of-the-art methods.
@@ II-A Computer vision applications of SNNs
The exploration of SNNs for object segmentation and image classification has ignited widespread interest in the field, as the SNNs replicate the biological neuronal system to provide efficient and effective solutions for this task. Initial works in object segmentation leaned on feed-forward SNN architectures, utilizing the spike-timing-dependent plasticity (STDP) learning rule for unsupervised learning of features [Deng2021STDPClassification]. The Backpropagation Through Time (BPTT) method introduced in [Kim2022BeyondSegmentation] for training SNNs, offers a more structured and supervised learning approach that significantly enhances feature extraction and classification accuracy, yet it is limited by its high computational complexity, substantial memory requirements, susceptibility to vanishing and exploding gradients, challenges in capturing long-term dependencies, and inefficiency in real-time learning applications. The introduction of Convolutional Spiking Neural Networks (CSNNs) marked a significant development in image classification. By incorporating convolutional layers into the network, these models could better handle the spatial complexities inherent in images, significantly enhancing segmentation performance. Concurrently, they maintained the temporal dynamics native to SNNs, enabling efficient processing of streaming data [Li2023ANetworks]. Recently, there has been a surge in the application of recurrent spiking neural networks (RSNNs) to image classification. These networks leverage feedback connections, enhancing the network’s ability to capture temporal dependencies in the data. Furthermore, their capacity to retain information over extended periods makes them particularly suited to segmentation tasks involving sequential or video data [Bohnstingl2022Biologically-inspiredHardware].
The memory-based methods such as those utilizing LSTM and RNN principles, highlighted by [bellec2018long] and [bellec2020solution], offer advancements yet face limitations due to their application across entire networks, constraining individual neuron-based threshold modeling. Similarly, Spiking Neural Networks (SNNs) have been applied to tasks like segmentation and object detection with promising results as demonstrated by [zhang2023accurate] and [zhang2023automotive]; however, these models often draw from Artificial Neural Network (ANN) principles, potentially veering away from the bio-plausibility sought to closely mimic biological brain functions. These developments underscore the challenges in achieving nuanced modeling and bio-inspired design within neural network research, highlighting the need for approaches that balance technical efficacy with the intricacies of biological mechanisms. Generative networks [leng2018spiking] in event-based vision struggle due to the asynchronous and unpredictable nature of the data, leading to performance issues in scenarios with sporadic events. This challenge highlights the complexity of creating bio-plausible models that maintain stability and accuracy in the face of such data variability.
@@ II-B Dynamic Thresholding in SNNs
Research on SNNs has delved deeply into dynamic thresholding methods to bolster adaptability. For instance, while predictive coding emphasizes anticipating input patterns, it overlooks the spike timing significance, an aspect STDP-based adjustments address [Rao1999PredictiveEffects], [Makhlooghpour2016HighModel]. The holistic approach of network activity regulation contrasts with the inherent refractory period modulation, focusing on preventing rapid neuron firing [WangRegulationModel]. Furthermore, spiking models like SRM and LIF have advanced neuron dynamics representations [Gerstner1995TimeModels., Gerstner2002SpikingPlasticity]. However, gaps remain, especially in dynamic threshold modulation. [Hao2018ARuleb] introduced a method utilizing a scaling factor to adjust the threshold in response to increased input. Yet, since it was tailored specifically for incremental spike membrane potential, it faltered when faced with variations in membrane decays. On the other hand, [Shaban2021AnImplementation.] adopted exponential functions to dictate decay. This approach, however, resulted in fluctuating thresholding levels, failing to ensure a smooth threshold curve. Meanwhile, [Kim2021SpikingUpdate.] ventured into a firing counter-based threshold adjustment. While innovative, it necessitated meticulous fine-tuning and exhaustive experimentation, leaving the determination of an optimal count mired in ambiguity.
Previous research [naeini2022event] may not addressed all the dynamics in the input signal, especially in highly complex event-based vision signals. In the event-based vision, SNNs face unique challenges: Neurons within the same layer exhibit varying event counts; some neurons experience more events than others. Consequently, applying uniformly the BDETT method [Ding2022BiologicallyNetworks], which utilizes average membrane potential of the neurons in a layer to predict the threshold of all the neurons in that layer, fails to reflect the individual dynamics of each neuron. This paper addresses notable research gaps. No prior studies have demonstrated that a bioinspired dynamic threshold can stabilize spike firing in semantic segmentation tasks. Additionally, these studies typically employ simplistic datasets such as NMNIST [Orchard2015ConvertingSaccades] and rarely evaluate system robustness under challenging conditions like motion blur or low light, vital for assessing the system’s homeostasis [Aydin2023APerception].
@@ III-A1 Event-based cameras
Event-based vision cameras are designed to register variations in logarithmic light intensities by recording individual pixel-level changes, referred to as events. These events formulate a continuous asynchronous event stream, which can be mathematically articulated as a series of ordered tuples. A tuple corresponds to an event $i$ and includes its spatial coordinates $(x_{i},y_{i})$ in the image frame, timestamps $t_{i}$, and polarity value $z_{i}$ [Lichtsteiner2008ASensor]:
@@ III-A2 Leaky integrate-and-fire
The LIF model, as shown in Fig. 1, is characterized by its linear integration of inputs until a threshold $\Theta(t)$ is reached, causing its firing [Gerstner2002SpikingModels]. Mathematically, the sub-threshold dynamics of the membrane potential $V(t)$ of an LIF neuron can be described by the following differential equation:
@@
where $\tau_{m}$ is the membrane time constant that dictates the speed at which the neuron responds to inputs, $E_{L}$ is the leak potential or resting potential representing the equilibrium potential of the neuron in the absence of any input, $R_{m}$ is the membrane resistance, and $I(t)$ represents the total synaptic input to the neuron that might come from other neurons or external sources. When the membrane potential $V(t)$ reaches the threshold $\Theta(t)$, the neuron emits a spike and then resets its potential to a reset value $V_{\text{reset}}$, typically below $\Theta(t)$. Additionally, after spiking, the neuron enters a refractory period for a duration $T_{\text{ref}}$ during which it cannot fire again. The ”leaky” aspect of the LIF model comes from the term $(V(t)-E_{L})$ in the differential equation. This term ensures that in the absence of any input, the neuron’s potential will gradually revert to the leak potential $E_{L}$.
@@ III-A3 Spike response model
In the SRM [Gerstner1995TimeModels.], incoming spike trains $s_{i}(t)$ are initially transformed into a spike response signal by convolving them with a spike response kernel $\varepsilon(\cdot)$, denoted as $(\varepsilon\ast s_{i})(t)$. Subsequently, this spike response signal is scaled by a synaptic weight $w_{i}$. The depolarization process involves summing up these scaled spike response signals across all incoming spikes, mathematically represented as $\Sigma_{i}w_{i}(\varepsilon\ast s_{i})(t)$. The SRM also incorporates a mechanism for hyperpolarization when a spike $s(t)$ is triggered, characterized by a refractory potential $(\zeta\ast s)(t)$, where $\zeta(\cdot)$ serves as the refractory kernel. This dual-component approach enables SRM to capture both excitatory and inhibitory dynamics in neuronal spike activity.
@@ III-A4 Asynchronous events to spike conversion
Traditionally, rate coding and temporal coding [Kim2022RateNetworks], [Lehmann2021SpikingBiCMOS] are the popular methods to convert frames, from a traditional CMOS, into spiking trains. In this study, similar to [Abreu2023FlowHardware], each event is treated as a spike to avoid the conversation of the input signal into a spike train. This approach essentially uses a form of temporal coding, because the information is conveyed through the timing of the spikes. We have chosen to drop the polarity aspect of the events due to its high sensitivity, which can introduce noise and complicate the signal processing. Mathematically, E(t) is the event stream, $e=(x,y,t)$ is an individual event with coordinates $x$, y and timestamp t, and $S_{i}(t)$ is the spike train for the i-th neuron:
@@
Here, $\delta(t-e_{t})$ is the Dirac delta function, which represents a ”spike” at time $t$.
@@ III-B Asynchronous Bioplausible Neuron
The equation models the dynamic adjustment of the neuron’s firing threshold at the next time step $t+1$ based on various factors. In this equation, $\Theta(t)$ is the existing threshold at time $t$, and $k_{1},k_{2},k_{3}$ are weights that determine the relative influence of each term. The term $k_{1}\cdot MG_{i}$ increases the threshold when the membrane potential is rapidly changing, preventing the neuron from firing too easily. Conversely, the term $k_{2}\cdot TRG_{i}$ decreases the threshold if there has been a historical trend of rapid threshold increases. Lastly, the term $k_{3}\cdot SE_{i}$ takes into account the neuron’s recent spiking activity. Specifically, if the neuron is overly active, this term will make it more difficult for the neuron to fire in the future, thereby ensuring efficiency. Overall, this equation encapsulates a multi-faceted approach to dynamically adapting the neuron’s firing threshold.
@@ III-B1 Membrane Gradient (MG)
The rate of change of the membrane potential is a critical parameter for predicting a dynamic threshold in SNNs. It serves as an indicator of the neuron’s responsiveness to incoming spikes and plays an essential role in governing the neuron’s firing behavior. Positive correlations between dynamic thresholds and membrane potentials have been observed in several areas of diverse biological nervous systems. In [Fontaine2014Spike-thresholdVivo.], the average membrane potential is correlated with an average threshold to predict the dynamic threshold. Instead, in this work we correlate the rate of change of membrane potential with a threshold. A high rate of change often signifies that the neuron is receiving a burst of spikes, necessitating a higher threshold to maintain neuronal homeostasis. Conversely, a lower rate of change could indicate that the neuron is less active, allowing for a lower dynamic threshold to be sensitive to sparse inputs. Incorporating the rate of change of the membrane potential into the calculation of a dynamic threshold ensures that the SNNs can adapt to varying synaptic activities, thereby emulating the adaptability and energy efficiency observed in biological neurons.
The BDETT method presented in [Ding2022BiologicallyNetworks] calculates the dynamic threshold using the average of membrane potentials and thresholds. In a layer with varied neuronal activity, not all neurons are always active. Including inactive neurons in this calculation can pull the average value lower. Applying this reduced average threshold to highly active neurons may result in excessive spiking, leading to overly aggressive neuronal activity. In contrast, in this work, we adopt the neuron-specific derivative term, denoted as $\eta\cdot\frac{dv_{i}(t)}{dt}$, which captures the rate of change of membrane potential of neuron $i$ at time $t$, denoted by $v_{i}(t)$. To numerically compute the derivative, a finite difference approximation is employed. This involves the subtraction of the membrane potential at the previous timestep from the current potential, followed by division by the timestep size $\Delta t$. The scaling factor $\eta$, which is set to a value less than 1, reduces the rate of change to ensure that the rate of change of the threshold remains less than the rate of change of the potential. The explicit equation for the derivative term is given as: